Compute the DTFTs for the following signals. In a. x[n] = (3) u[-n – 1] b. x[n] = 2” sin (n)u[-n]
Answer #1

Q: Soin in XuJ= (4) . U[-noi ras DTFT formula -Jww. DTFT {REng} = Žaluj-é +00 ? N =D Here a[u]=ļ (1/2) ne-l pluta otherwise .<by x[u]=%. Sin (%) u[n] DTFT formula و مالی stu].e DTFT{Xn]}= m-do Here, ~ sin ( /m) ; nxo X[u]= . other wise e * sind 409[ (ده به پW) 1 ( = •e = 6 - (1 ) X tup to + ا e 3 : رما از مد | Now, سے = (سانہ )X ( 4 ) ) نے ع--| (س) ز ار ع-2 -1 ТЕТ 1.86]